At General Hospital, nurses are given performance evaluations to determine eligibility for merit pay raises. The supervisor rates the nurses on a scale of 1 to 10 (10 being the highest rating) for several activities: promptness, record keeping, appearance, and bedside manner with patients. Then an average is determined by giving a weight of 4 for promptness, 2 for record keeping, 3 for appearance, and 1 for bedside manner with patients. What is the average rating for a nurse with ratings of 5 for promptness, 2 for record keeping, 10 for appearance, and 2 for bedside manner? (Enter your answer to one decimal place.)
Answer #1
type rating(x) weight(w) w*x
promptness 5 4 20
record keeping 2 2 4
appearance 10 3 30
bedside manner 2 1 2
total 10 56

  average rating \sum xi*wi/\sumwi =56/10 =5.6